An eastbound train left at noon. Four hours later, a westbound train left the same station. At eight o'clock the two trains were 400 miles apart. Find the rate of each train if the westbound train traveled 10 miles per hours faster than the eastbound train.
Answer by stanbon(75887)

An eastbound train left at noon. Four hours later, a westbound train left the same station. At eight o'clock the two trains were 400 miles apart. Find the rate of each train if the westbound train traveled 10 miles per hours faster than the eastbound train.
-------
Eastbound DATA:
time = 8 hrs ; rate = x mph ; distance = 8x miles
-----
Westbound DATA:
time = 4 hrs ; rate = x+10 ; distance = 4x+40 miles
------
distance - distance = 400 miles
8x - (4x+40) = 400
4x = 360
x = 90 mph (rate of Eastbound)
x+10 = 100 mph (rate of Westbound)
====================================
Cheers,
Stan H.
==================